The Atlantic Company sells a product with a break-even point of 3,538 sales units. The variable cost is $69 per unit, and fixed costs are $106,140.

Determine the unit sales price. Round answer to nearest whole number.
$

Determine the break-even point in sales units if the company desires a target profit of $57,420. Round answer to the nearest whole number.
units
Answer #1

Answers :

A) Unit Selling Price: $ 99

B) BEP To get he Desired Profit of $57,420 = 5,452 Units

Unit sales Price = ( Fixed cost + Varaible Cost + Profir(loss))/Units = ( 106,140+(69*3538)+0)/3538 = $ 99

Units to get the Desired Profit of $ 57,420 = ( Fixed cost + Desired Profit) / CM per unit

= ($ 106,140+57,420) /( 99-69) = 5,452 Units